### v2

# Evidence for itinerant electron-local moment interaction in Li-doped -MnTe Tingjun Zhang

Abstract
We use inelastic neutron scattering (INS) and angle-resolved photoemission spectroscopy (ARPES) to study the impact of Li doping on the semiconducting altermagnet -MnTe.
Introducing Li results in a spin reorientation from in-plane to out-of-plane
direction and increases the density of itinerant carriers.
While our ARPES measurements do not indicate any notable doping-induced changes in the electronic band structure or the magnitude of the altermagnetic band splitting, our INS measurements reveal an abrupt carrier-induced decrease in the spin wave lifetime near the zone boundary at high energies. This finding is consistent with a new magnon decay channel driven by doping-induced subtle changes in the band structure and enhanced interactions between Mn2+ local moments and itinerant electrons. By extracting the local dynamic susceptibility from INS spectra and applying the total moment sum rule, we find that both undoped and Li-doped -MnTe exhibit the full expected Mn2+ local moment of with . These findings suggest that -MnTe hosts robust local-moment altermagnetism which shows a breakdown at high energies upon addition of itinerant carriers, highlighting the importance of carrier-spin coupling in magneto-transport and spin dynamic properties of altermagnets even in the dilute-carrier limit.
In conventional local-moment ferromagnets and antiferromagnets with dipolar magnetic exchange interactions,
the collective motion of localized moments in the magnetic ordered state is spin waves (magnons) characterized by linear spin wave theories (LSWT) that ignore all terms of order higher than quadratic [32, 33, 50, 7].
Their dispersion determines the nearest-neighbor (NN), next-nearest-neighbor (NNN), etc. exchange couplings, and, therefore, the magnetic ordering temperature – regardless of the ordered moment direction controlled by the spin-orbit coupling (SOC)
[7].
Since magnon-magnon interactions are negligible [74], the magnon lifetime as measured by the energy width of spin wave spectra should be infinite and only limited by the instrumental resolution, as seen in ferromagnetic EuO [17] and collinear antiferromagnetic Rb2MnF4 [5].
For classical spin systems (i.e.,) without strong quantum fluctuations, geometric frustration, or SOC,
magnon decay processes are generally not expected in insulating magnets [74].
Figure 1: (a,b) Parallel view of the lattice and magnetic structures of MnTe and Li-MnTe, respectively, highlighting magnetic moment directions and exchange couplings. The dashed lines indicate the altermagnetic exchange interactions and . The Li occupancy at Te-site defects is exaggerated.
(c,d) Schematics of magnetic domain configurations MnTe and Li-MnTe, respectively.
(e) Magnetization of Li-MnTe versus temperature .
(f) Longitudinal resistivity versus for MnTe and Li-MnTe.
(g) Transverse resistivity of Li-MnTe.
(h) Carrier density and mobility of MnTe and Li-MnTe.
(i,j) Calculated spin waves of MnTe for three domains within (j) and outside (k) the nodal planes.
The calculations use exchange parameters from Ref. [45].
The inset in (k) displays the intensity contributions from the three domains at the position of altermagnetic splitting. The solid and dashed lines denote different branches of the spin waves, and the vertical dashed line separates the splitting dominated by single-ion anisotropy () and altermagnetism ().
(k,l) Calculated spin waves of Li-MnTe, with the single-ion anisotropy replaced by the value reported in Ref. [73].
In metallic magnets, spin-wave excitations can decay through their coupling to conduction electrons. In strongly itinerant systems, magnons may enter the Stoner particle-hole continuum at elevated energies, where their lifetime is rapidly reduced by Landau damping arising from decay into electron-hole spin-flip excitations [65, 66, 60, 71]. In local-moment metals, the magnetic dispersion is primarily governed by the dipolar exchange interactions between well-defined local moments, while low-energy scattering of conduction electrons can produce Korringa-type damping that is most clearly manifested at long-wavelengths. The hydrodynamic form of such relaxation processes is commonly described phenomenologically within the Landau-Lifshitz-Gilbert framework [43, 26], which captures viscous damping of collective spin precession but does not by itself specify the underlying microscopic mechanism. In many three-dimensional metallic ferromagnets, this intrinsic damping remains relatively weak over a broad energy range; for example, in elemental -Fe spin waves are observed to remain essentially resolution-limited up to energies approaching the onset of the Stoner continuum near 200 meV [46, 52, 9]. Consequently, measurable spin-wave broadening in metallic magnets provides a sensitive probe of the degree of itinerancy and of the details of carrier-spin coupling encoded in the electronic structure near the Fermi level.
In the semiconducting altermagnet -MnTe (hereafter, MnTe), the opposite magnetic sublattices of Mn2+ are connected by symmetry operations that are not inversion () or translation () symmetric due to the specific local crystal environment [Fig. 1(a)]. This leads to alternating spin splitting of electronic energy bands, arising from exchange coupling rather than SOC [61, 62, 63]. This is in sharp contrast to most collinear antiferromagnets, where the combination of time-reversal symmetry () and or symmetries between opposite magnetic sublattices enforces the degeneracy of electronic bands throughout the entire Brillouin zone (BZ) in the limit of zero SOC [7]. Indeed, several phenomena commonly associated with the breaking of -symmetry – spin-split electronic bands [18], anomalous Hall effects (AHE) [27, 36, 6, 44, 64],
and magnetic circular dichroism [31] – have been experimentally realized in altermagnetic materials [4, 25].
In addition to momentum-dependent electronic energy band splitting, altermagnetic symmetry also gives rise to a momentum-dependent energy splitting of chiral magnons which is non-relativistic in origin and holds even in the limit of symmetric exchange interactions [45, 59, 67, 57, 24].
An attractive aspect of MnTe in the context of spintronics is the tunability of its altermagnetic ground state, which can be modified by
chemical substitution [48, 73],
uniaxial strain [6, 44, 64],
and external applied fields [40, 38, 2, 28, 19]. MnTe adopts a layered -type magnetic order below K [41, 20, 69, 39, 68], where Mn2+ moments are oriented in-plane along the NNN Mn-Mn direction [Fig. 1(a,c)] [44].
The hexagonal in-plane symmetry enforces the coexistence of three degenerate magnetic domains, each rotated 120∘ from the other [Fig. 1(c)], resulting in a magnetic response which is a superposition of all three.
The anisotropy of these Mn2+ moments can be strongly modified from in-plane to out-of-plane by less than 1% Li doping, which also introduces charge carriers into the system [Fig. 1(b,d)] [73].
This spin reorientation removes the three-fold degeneracy in the magnetic channel, and effectively provides another pathway to achieving a single domain response without modifying the symmetries required for altermagnetism [Fig. 1(d)] [61, 62, 63].
Figure 2: (a,b) INS spectra acquired at base temperature ( K) in HF mode with an incident energy of meV for Li-MnTe and MnTe, shown along high-symmetry directions within the nodal plane.
The black solid and dashed curves represent calculated spin wave branches, which are degenerate in the Li-doped sample.
(c) FWHM values obtained from Gaussian fits of the INS spectra in the nodal planes of the BZ for pure and Li-MnTe. Solid and dashed curves represent the energy-dependent instrumental resolutions from Fig. S1.
(d) FWHM values obtained by subtracting the FWHM of pristine MnTe from that of Li-MnTe. Vertical error bars represent 1 standard deviations and horizontal error bars indicate the integration range along the corresponding path.
(e) Energy-dependent line cuts at selected points along the -M path. The fitted FWHM values are listed within each sub-panel.
(f) Schematic of the BZ, where nodal planes are shaded in gray and the -A–L–M plane is shaded in red.
Filled symbols use Li-MnTe HF data and open symbols use Li-MnTe HR data in each panel.
Understanding the influence of charge carriers on the prototypical altermagnetic ground state of MnTe and on its collective excitations is important. The AHE, for example, has been shown to depend sensitively on sample-dependent variations in the charge gap, indicating that a narrow window of in-gap impurity states is required for its observation in MnTe [44]. At the microscopic level, the chiral splitting of the magnon branches originates from the inequivalence of the exchange pathways and which, despite having equal bond lengths, are distinguished by their distinct superexchange routes through Te orbitals [Fig. 1(a)] [45]. In Li-doped MnTe (hereafter, Li-MnTe), the extent to which the introduction of charge carriers and the associated substitutional disorder affect the spin dynamics that underpin altermagnetic functionality remains unresolved. To address this, we investigate the effects of Li doping on the magnetic excitations and electronic structure of Li-MnTe using single-crystal inelastic neutron scattering (INS) and angle-resolved photoemission spectroscopy (ARPES), respectively. We find that Li doping produces several notable changes, particularly a pronounced magnon damping near the zone boundary above meV, pointing to an increased role of interactions between itinerant electrons and Mn2+ local moments which are essential for understanding the emergence of the AHE [44] as well as the spin dynamics in this system.
First, we attempt to determine the position and occupancy of Li by refining several structural models against single crystal neutron diffraction data.
Our results indicate an approximately 4% Te-site deficiency in both MnTe and Li-MnTe (Tables S5–S11).
In Li-MnTe, allowing Li to occupy the Te site yields a stable refined Li occupancy of 3% at both 90 K and 340 K, whereas placing Li on the site or on tested interstitial positions drives the Li occupancy to values that are effectively zero and does not improve the agreement factors (Table S3).
While diffraction alone is not sensitive to Li incorporated on uncorrelated interstitial sites, these results suggest that Li is preferentially accommodated on the Te site at a level of 3%.
The magnetization of the 3% Li-doped sample is consistent with a Néel vector along the axis [Fig. 1(e)] [73], while the resistivity drops by two orders of magnitude relative to MnTe [Fig. 1(f)].
Notably, remains unchanged at this doping level, reminiscent of the robust antiferromagnetic ground state observed in K-doped BaMn2As2 [42].
This suggests that the effective exchange energy scale between Mn2+ local moments remains similarly unchanged.
The extracted carrier density and positive Hall coefficient indicate that holes are the dominant carrier in both systems, and Li doping enhances their contribution to conduction [Fig. 1(g,h)].
To understand the impact of Li-doping on the spin waves of Li-MnTe,
we conducted time-of-flight INS measurements at the SEQUOIA spectrometer at the Spallation Neutron Source for pristine MnTe and at the MAPS Spectrometer at the ISIS Neutron Source for Li-MnTe. Measurements were performed on a single piece of single crystal with mosaic less than 1 degree, using nearly identical instrumental resolutions (Fig. S1) and normalized to absolute units in order to ensure a quantitative comparison of the experimental spectra [72].
In pristine MnTe, chiral magnon splitting appears in the -A-L-M plane, highlighted in red in Fig. 2f, while spin waves remain degenerate in the nodal planes shaded in gray [45].
We first examine spin waves within the nodal planes [Fig. 2 (a,b)], which reveal three differences induced by Li doping:
(1) the spectra evolve from pseudo-gapless to having a 6 meV spin gap, indicative of the change in single-ion anisotropy,
(2) the overall bandwidth of the excitations is reduced from 35 meV to 32 meV, signaling a 9% reduction in the effective exchange energy scale, and
(3) whereas spin waves in pristine MnTe are resolution-limited throughout the BZ, the linewidth in Li-MnTe is notably larger near the zone boundary [Fig. 2(c-e)].
Whereas (2) can be understood as the addition of indirect exchange via Ruderman-Kittel-Kasuya-Yosida (RKKY)-type interactions mediated by itinerant carriers, the appearance of magnon damping implied by (3) may have several possible physical origins in metallic magnets, as previously discussed.
To distinguish between these scenarios, we performed energy-dependent Gaussian profile fits at constant to compare the energy full width at half-maximum (FWHM) between MnTe and Li-MnTe in high-flux (HF) mode, where the instrumental resolutions are essentially identical (Fig. S1). We also carried out high-resolution (HR) measurements for Li-MnTe, in order to differentiate between intrinsic and extrinsic broadening effects (Figs. S2, S3). While the linewidth at the point is dominated by the presence of a nearby optical mode [Fig. 1(i,k)], by moving slightly away from the point one can draw a quantitative comparison between the FWHM of Li-MnTe and MnTe. Taking the MnTe FWHM to represent the intrinsic linewidth, the difference profile (FWHM) between the two HF datasets demonstrates that the linewidth is nearly identical between the two samples away from the BZ boundary [Fig. 2(d)], with the exception of an artifact at intermediate (e.g., halfway along -M) which manifests as a hump in the HF dataset but is less pronounced in the HR dataset. In contrast, near the zone boundary we find a notable increase in the FWHM of Li-MnTe in both HF and HR modes, strongly suggesting that this is an intrinsic effect induced by Li doping. Examining the energy-dependence of this broadening reveals that its onset is limited to energies above meV, with slightly different onset energies at different (Fig. S4).
Figure 3: (a,b) INS spectra along and (c,d) along the direction for MnTe and Li-MnTe.
(e) Schematic of the -A-L-M plane of the BZ, where splitting is expected to occur in MnTe.
Green and purple lines highlight the and directions, respectively.
Arrows indicate the direction of constant- line cuts.
(f) Comparison of line cuts for the dispersion along the and
(g) directions.
Points represent the measured data, and error bars represent 1 uncertainties.
Solid and dashed curves represent fits to LSWT models where and are nonzero and zero, respectively.
For Li-MnTe, the nonzero values of and were fixed to the refined values from MnTe.
Data for Li-MnTe was shifted vertically by 0.075 units for clarity.
Such an abrupt reduction in the spin-wave lifetime is inconsistent with the smooth energy dependence expected for disorder-dominated damping [35, 8, 12]. In the dilute-defect limit, the magnon self-energy is governed by the single-impurity -matrix and scales linearly with defect concentration [35, 8]. For strong point-like scatterers, the resulting linewidth is expected to follow , where is the magnon energy, implying
that the linewidth increase should be proportional to
impurity level ; and should
gradually increase with rather than a sharp onset [12]. Comparable behavior is obtained in supercell and coherent-potential simulations of disordered Heisenberg magnets [10]. The abrupt broadening also differs qualitatively from the damping attributed to dynamical spin-lattice coupling in ferromagnet CrGeTe3 [11], which produces a more gradual energy evolution. Since spin-lattice coupling is not expected to change substantially upon light Li substitution, the pronounced high-energy damping in Li-doped MnTe instead points to an additional decay channel associated with increased itinerant character and interactions between itinerant electrons and local moments. A related finite linewidth broadening has been reported in low-carrier-density Mn-based metallic magnets such as YbMnBi2 [55] and YbMnSb2 [34].
As a measure of the modifications to the spin Hamiltonian within a local-moment framework, we fit the spin wave spectra of both compounds using LSWT based on a Heisenberg Hamiltonian with a single-ion anisotropy term (Eqn. S1). A positive (negative) corresponds to the easy-plane (easy-axis) anisotropy in the undoped (doped) system. Compared to MnTe, the NN exchange coupling in Li-MnTe is reduced by 8%, consistent with the spin wave bandwidth reduction. To quantify the possible impurity-induced exchange-disorder mechanism, we modeled the measured energy dependence of the spectral weight at each momentum transfer [] as a Gaussian profile with a finite HWHMFWHM/2. We then constructed three independent LSWT models for Li-MnTe based on distinct input dispersions along the BZ-boundary segments M-L and L-A, where the linewidth broadening is most pronounced. The reference (“center”) model employed the dispersion defined by the peak energies obtained directly from Gaussian fits to the experimental spectra. Two additional models were generated by systematically shifting these peak energies at each along these boundary paths by the corresponding linewidth, yielding upper and lower bounding dispersions defined by and , respectively. These three cases are denoted as “upper”, “center”, and “lower” in Table S1.
Among the exchange parameters, the further-neighbor interactions and the anisotropy term were found to be relatively insensitive to this spectral weight distribution. In contrast, the term exhibited a distribution range of meV, corresponding to a HWHM of about 1.5 meV in the spectral weight at the BZ boundary or about 10% of . Therefore, in Li-MnTe is 3 times larger than , again inconsistent with impurity-induced effect [35, 8, 12].
We now inspect the spin wave spectra on the -A-L-M plane of the BZ, where splitting is expected in MnTe [Fig. 3(e)] [45]. Figure 3(a-d) show a comparison of spin wave spectra along the and directions for both systems.
While splitting is clearly observed in MnTe [45], the broadened linewidth in Li-MnTe makes the splitting more challenging to resolve.
This is also reflected in constant energy cuts [Fig. 4(a-c)], where inspection of the -dependence of the ring-like spectral weight reveals an alternating bright/dim pattern around the ring, which is a direct result of the splitting in MnTe. We note that in Li-MnTe a weak alternation is evident at 34.5 meV, which is consistent with a persistent splitting [Fig. 4(c)].
To more closely investigate the presence of magnon splitting in the two systems, we performed constant- cuts near the expected splitting position in - space [Fig. 3(e-g)].
While two peaks are resolvable in the line cuts for MnTe, the spectra for Li-MnTe are evidently broader and only a single, broad feature is resolvable for most values.
Figure 4: (a-c) Constant-energy maps of the INS spectra in the plane for pure (left) and Li-MnTe (right) between 33 meV and 35 meV. The BZ is outlined by a thin grey line. The observed splitting in pure MnTe is indicated by red arrows. (d) Energy dependence of the local dynamic susceptibility extracted from INS data, with effective moments obtained from the total moment sum rule. (e) ARPES 100 meV integrated constant energy contour of Li-MnTe centered around the Fermi level, showing the first and second electronic BZ. (f) Second energy derivative high-symmetry cut along K--K with altermagnetic splitting of meV indicated by red bars.
To understand whether the observed dispersion and lineshape for Li-MnTe can be explained by a persistent splitting which is simply broadened by the addition of itinerant carriers, we fit the data to two distinct LSWT models with refined , , : (1) with and zero, where splitting is necessarily absent; (2) with non-zero difference fixed at the value refined for MnTe. and were allowed to refine under this constraint [Fig. 3(f,g)].
The former scenario, which is identical to the “Li-doped MnTe center” presented earlier, underestimates both the spin wave energy and the linewidth near the expected splitting (Table S1).
For the model with , we find improved agreement with experimental energies, further suggesting that the splitting in MnTe may survive Li-doping.
While Li-doping strongly modifies the spin anisotropy and tunes the electronic carrier concentration, altermagnetism is, a priori, expected to remain robust to a first-order approximation as its framework holds valid for both insulating and metallic systems.
To verify this, we performed ARPES measurements on Li-MnTe to interrogate the behavior of the electronic band splitting in the presence of Li dopants. As with MnTe, the Fermi surface at consists of six petal-like pockets around [Fig. 4(e)] [37, 51]. Along -K, where altermagnetic splitting from SOC has previously been predicted and observed in MnTe, we find splitting of approximately the same magnitude (300 meV) [37, 51, 18], verifying the robust nature of altermagnetism in 3% Li-doped MnTe [Fig. 4(f)]. Within the scope of these experiments, no obvious changes in the band structure near the Fermi level were observed when compared to MnTe. Thus, the observed magnon damping must arise from subtle changes in the electronic structure that impact the nature of the magnetic exchange and spin dynamics via coupling between itinerant electrons and local moments through RKKY (indirect exchange) interactions and magnon-electron scattering processes, respectively.
The fact that small amounts of Li doping can dramatically affect the spin gap and the spin wave bandwidth is also supportive of the important role of itinerant electrons in this system.
Further, the sensitive dependence of the AHE on the charge gap in undoped MnTe [44] underscores the notion that the physics of this system can largely be described as a network of local moments that interact with itinerant charge carriers. We further test this notion by calculating the local dynamic susceptibility from the total spectral weight of our INS measurements [Fig. 4(d)], which, when integrated across all energy transfers, yields a quantitative measure of the local dynamic moment [72, 16]. Our analysis yields 1.53(1) [1.84(6) ] for pristine MnTe [Li-MnTe]. When combined with the corresponding static ordered moment (Eqn. S2), this indicates a total effective moment of 5.88(4) [6.1(2) ].
The small but finite magnon broadening is not expected to remove spectral weight, and none of the spectral weight is pushed to energies outside the measurement range. Thus, these values are close to the expected for () Mn2+ ions, therefore strongly suggesting a local moment picture for altermagnetism in both compounds despite partially itinerant character at high energies upon 3% Li doping.
In summary, magnon damping is observed in Li-MnTe along multiple high-symmetry paths near the BZ boundary at high energy, suggesting a mechanism based in subtle electronic structure changes with a characteristic energy distinct from the strongly momentum-dependent altermagnetic splitting.
Additional chirality-dependent magnon decay processes have been proposed in altermagnets [15, 13, 21], although whether these mechanisms can be directly observed by INS remains an open question.
This finding underscores the important role of itinerant electron-local moment interactions, which must be considered to understand the physics of altermagnetic candidate materials and their potential for applications in next-generation spintronic devices.
We thank Kirill Belashchenko, Erick Lawrence, and Yichen Zhang for helpful discussions.
The single-crystal synthesis and neutron scattering work at Rice were supported by the U.S. DOE, BES under Grant Nos. DE-SC0026179 (P.D., M.Y., and E.M.) and DE-SC0012311 (P.D.). Part of the materials characterization efforts at Rice is supported by the Robert A. Welch Foundation Grant No. C-1839 (P.D.). X.W. acknowledges research sponsored by the Laboratory Directed Research and Development Program of ORNL, managed by UT-Battelle, LLC, for the U.S. DOE. Part of the materials characterization efforts at University of Houston is supported by the T. L. L. Temple Foundation; the John J and Rebecca Moores Endowment; the State of Texas through the Texas Center for Superconductivity at the University of Houston; and the Robert A. Welch Foundation (00730-5021-H0452-B0001-G0512489). This research was supported in part by an appointment to the ORNL GRO Program, sponsored by the U.S. DOE and administered by the Oak Ridge Institute for Science and Education.
A portion of this research used resources at the Spallation Neutron Source, a DOE Office of Science User Facility operated by ORNL. The beam times were allocated to TOPAZ and SEQUOIA on Proposals No. IPTS-35052, IPTS-33830 and IPTS-34120. The experiment at the ISIS Neutron and Muon Source was supported by beam time allocations RB2510033 and RB2510079 from the Science and Technology Facilities Council. The data supporting this manuscript are available from the authors upon reasonable request.
[2]O. J. Amin, A. Dal Din, E. Golias, Y. Niu, A. Zakharov, S. C. Fromage, C. J. B. Fields, S. L. Heywood, R. B. Cousins, F. Maccherozzi, J. Krempaský, J. H. Dil, D. Kriegner, B. Kiraly, R. P. Campion, A. W. Rushforth, K. W. Edmonds, S. S. Dhesi, L. Šmejkal, T. Jungwirth, and P. Wadley (2024-12)Nanoscale imaging and control of altermagnetism in MnTe.
Nature636 (8042), pp. 348–353.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[3]O. Arnold, J. Bilheux, J. Borreguero, A. Buts, S. I. Campbell, L. Chapon, M. Doucet, N. Draper, R. F. Leal, M. Gigg, et al. (2014)Mantid—Data analysis and visualization package for neutron scattering and SR experiments.
Nucl. Instrum. Methods Phys. Res. A764, pp. 156–166.
Cited by: §I.3.
[4]A. Barman, G. Gubbiotti, S. Ladak, A. O. Adeyeye, M. Krawczyk, J. Gräfe, C. Adelmann, S. Cotofana, A. Naeemi, V. I. Vasyuchka, B. Hillebrands, S. A. Nikitov, H. Yu, D. Grundler, A. V. Sadovnikov, A. A. Grachev, S. E. Sheshukova, J. Duquesne, M. Marangolo, G. Csaba, W. Porod, V. E. Demidov, S. Urazhdin, S. O. Demokritov, E. Albisetti, D. Petti, R. Bertacco, H. Schultheiss, V. V. Kruglyak, V. D. Poimanov, S. Sahoo, J. Sinha, H. Yang, M. Münzenberg, T. Moriyama, S. Mizukami, P. Landeros, R. A. Gallardo, G. Carlotti, J. Kim, R. L. Stamps, R. E. Camley, B. Rana, Y. Otani, W. Yu, T. Yu, G. E. W. Bauer, C. Back, G. S. Uhrig, O. V. Dobrovolskiy, B. Budinska, H. Qin, S. van Dijken, A. V. Chumak, A. Khitun, D. E. Nikonov, I. A. Young, B. W. Zingsem, and M. Winklhofer (2021-08)The 2021 Magnonics Roadmap.
J. Phys. Condens. Matter33 (41), pp. 413001.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[5]S. P. Bayrakci, D. A. Tennant, Ph. Leininger, T. Keller, M. C. R. Gibson, S. D. Wilson, R. J. Birgeneau, and B. Keimer (2013-07)Lifetimes of Antiferromagnetic Magnons in Two and Three Dimensions: Experiment, Theory, and Numerics.
Phys. Rev. Lett.111, pp. 017204.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[11]L. Chen, C. Mao, J. Chung, M. B. Stone, A. I. Kolesnikov, X. Wang, N. Murai, B. Gao, O. Delaire, and P. Dai (2022-JUL 12)Anisotropic magnon damping by zero-temperature quantum fluctuations in ferromagnetic CrGeTe3.
Nat. Commun.13 (1), pp. 4037.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[14]L. Coates, H. Cao, B. C. Chakoumakos, M. D. Frontzek, C. Hoffmann, A. Y. Kovalevsky, Y. Liu, F. Meilleur, A. M. dos Santos, D. A. Myles, et al. (2018)A suite-level review of the neutron single-crystal diffraction instruments at oak ridge national laboratory.
Rev. Sci. Instrum.89 (9).
External Links: LinkCited by: §I.2.
[19]J. Dzian, P. Kubaščík, S. Tázlarů, M. Białek, M. Šindler, F. Le Mardelé, C. Kadlec, F. Kadlec, M. Gryglas-Borysiewicz, K. P. Kluczyk, A. Mycielski, P. Skupiński, J. Hejtmánek, R. Tesař, J. Železný, A.-L. Barra, C. Faugeras, J. Volný, K. Uhlířová, L. Nádvorník, M. Veis, K. Výborný, and M. Orlita (2025-07)Antiferromagnetic resonance in -MnTe.
Phys. Rev. B112 (2), pp. 024433.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[21]R. Eto, M. Gohlke, J. Sinova, M. Mochizuki, A. L. Chernyshev, and A. Mook (2025)Spontaneous magnon decays from nonrelativistic time-reversal symmetry breaking in altermagnets.
Phys. Rev. B112 (9), pp. 094442.
Cited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[22]R. Ewings, A. Buts, M. Le, J. van Duijn, I. Bustinduy, and T. Perring (2016)Horace: software for the analysis of data from single crystal spectroscopy experiments at time-of-flight neutron instruments.
Nucl. Instrum. Methods Phys. Res. A834, pp. 132–142.
Cited by: §I.3.
[23]R. Ewings, J. Stewart, T. Perring, R. Bewley, M. Le, D. Raspino, D. Pooley, G. Škoro, S. Waller, D. Zacek, et al. (2019)Upgrade to the MAPS neutron time-of-flight chopper spectrometer.
Rev. Sci. Instrum.90 (3).
Cited by: §I.3.
[25]B. Flebus, D. Grundler, B. Rana, Y. Otani, I. Barsukov, A. Barman, G. Gubbiotti, P. Landeros, J. Akerman, U. Ebels, P. Pirro, V. E. Demidov, K. Schultheiss, G. Csaba, Q. Wang, F. Ciubotaru, D. E. Nikonov, P. Che, R. Hertel, T. Ono, D. Afanasiev, J. Mentink, T. Rasing, B. Hillebrands, S. V. Kusminskiy, W. Zhang, C. R. Du, A. Finco, T. van der Sar, Y. K. Luo, Y. Shiota, J. Sklenar, T. Yu, and J. Rao (2024-06)The 2024 magnonics roadmap.
J. Phys. Condens. Matter36 (36), pp. 363501.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[27]R. D. Gonzalez Betancourt, J. Zubáč, R. Gonzalez-Hernandez, K. Geishendorf, Z. Šobáň, G. Springholz, K. Olejník, L. Šmejkal, J. Sinova, T. Jungwirth, S. T. B. Goennenwein, A. Thomas, H. Reichlová, J. Železný, and D. Kriegner (2023-01)Spontaneous Anomalous Hall Effect Arising from an Unconventional Compensated Magnetic Phase in a Semiconductor.
Phys. Rev. Lett.130 (3), pp. 036702.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[28]R. D. Gonzalez Betancourt, J. Zubáč, K. Geishendorf, P. Ritzinger, B. Růžičková, T. Kotte, J. Železný, K. Olejník, G. Springholz, B. Büchner, A. Thomas, K. Výborný, T. Jungwirth, H. Reichlová, and D. Kriegner (2024-08)Anisotropic magnetoresistance in altermagnetic MnTe.
npj Spintronics2 (1), pp. 45.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[29]G. E. Granroth, D. H. Vandergriff, and S. E. Nagler (2006)SEQUOIA: A fine resolution chopper spectrometer at the SNS.
Physica B385, pp. 1104–1106.
Cited by: §I.3.
[30]G. Granroth, A. Kolesnikov, T. Sherline, J. Clancy, K. Ross, J. Ruff, B. Gaulin, and S. Nagler (2010)SEQUOIA: a newly operating chopper spectrometer at the SNS.
In J. Phys.: Conf. Ser.,
Vol. 251, pp. 012058.
Cited by: §I.3.
[31]A. Hariki, A. Dal Din, O. J. Amin, T. Yamaguchi, A. Badura, D. Kriegner, K. W. Edmonds, R. P. Campion, P. Wadley, D. Backes, L. S. I. Veiga, S. S. Dhesi, G. Springholz, L. Šmejkal, K. Výborný, T. Jungwirth, and J. Kuneš (2024-04)X-Ray Magnetic Circular Dichroism in Altermagnetic -MnTe.
Phys. Rev. Lett.132 (17), pp. 176701.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[34]X. Hu, A. Sapkota, Z. Hu, A. T. Savici, A. I. Kolesnikov, J. M. Tranquada, C. Petrovic, and I. A. Zaliznyak (2023-05)Coupling of magnetism and Dirac fermions in ${\mathrm{YbMnSb}}_{2}$.
Phys. Rev. B107 (20), pp. L201117.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[36]K. P. Kluczyk, K. Gas, M. J. Grzybowski, P. Skupiński, M. A. Borysiewicz, T. Fąs, J. Suffczyński, J. Z. Domagala, K. Grasza, A. Mycielski, M. Baj, K. H. Ahn, K. Výborný, M. Sawicki, and M. Gryglas-Borysiewicz (2024-10)Coexistence of anomalous Hall effect and weak magnetization in a nominally collinear antiferromagnet MnTe.
Phys. Rev. B110 (15), pp. 155201.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[37]J. Krempaský, L. Šmejkal, S. D’Souza, H. M, S. G, K. Uhlířová, F. Alarab, P. Constantinou, V. Strocov, D. Usanov, W. Pudelko, R. González-Hernández, A. Birk Hellenes, Z. Jansa, H. Reichlová, Z. Šobáň, R. Gonzalez Betancourt, P. Wadley, J. Sinova, D. Kriegner, J. Minár, J. Dil, and T. Jungwirth (2024-02)Altermagnetic lifting of kramers spin degeneracy.
Nature626 (7999), pp. 517–522.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[38]D. Kriegner, H. Reichlova, J. Grenzer, W. Schmidt, E. Ressouche, J. Godinho, T. Wagner, S. Y. Martin, A. B. Shick, V. V. Volobuev, G. Springholz, V. Holý, J. Wunderlich, T. Jungwirth, and K. Výborný (2017-12)Magnetic anisotropy in antiferromagnetic hexagonal MnTe.
Phys. Rev. B96 (21), pp. 214418.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[39]D. Kriegner, H. Reichlova, J. Grenzer, W. Schmidt, E. Ressouche, J. Godinho, T. Wagner, S. Y. Martin, A. B. Shick, V. V. Volobuev, G. Springholz, V. Holý, J. Wunderlich, T. Jungwirth, and K. Výborný (2017)Magnetic anisotropy in antiferromagnetic hexagonal MnTe.
Phys. Rev. B96 (21), pp. 214418.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[40]D. Kriegner, K. Výborný, K. Olejník, H. Reichlová, V. Novák, X. Marti, J. Gazquez, V. Saidl, P. Němec, V. V. Volobuev, G. Springholz, V. Holý, and T. Jungwirth (2016-06)Multiple-stable anisotropic magnetoresistance memory in antiferromagnetic MnTe.
Nat. Commun.7 (1), pp. 11623.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[42]J. Lamsal, G. S. Tucker, T. W. Heitmann, A. Kreyssig, A. Jesche, A. Pandey, W. Tian, R. J. McQueeney, D. C. Johnston, and A. I. Goldman (2013)Persistence of local-moment antiferromagnetic order in Ba${}_{1\ensuremath{-}x}$K${}_{x}$Mn${}_{2}$As${}_{2}$.
Phys. Rev. B87 (14), pp. 144418.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[47]K. Momma and F. Izumi (2011)VESTA 3 for three-dimensional visualization of crystal, volumetric and morphology data.
J. Appl. Crystallogr.44 (6), pp. 1272–1276.
Cited by: §I.2.
[48]D. H. Moseley, K. M. Taddei, J. Yan, M. A. McGuire, S. Calder, M. M. H. Polash, D. Vashaee, X. Zhang, H. Zhao, D. S. Parker, R. S. Fishman, and R. P. Hermann (2022-01)Giant doping response of magnetic anisotropy in MnTe.
Phys. Rev. Mater.6 (1), pp. 014404.
External Links: DocumentCited by: §II.1,
Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[53]V. Petříček, L. Palatinus, J. Plášil, and M. Dušek (2023)Jana2020–a new version of the crystallographic computing system Jana.
Z. Kristallogr. Cryst. Mater.238 (7-8), pp. 271–282.
Cited by: §I.2.
[54]V. Reshniak, X. Wang, G. Zhang, S. Liu, and J. Yin (2024)Hierarchical bayesian approach for adaptive integration of bragg peaks in time-of-flight neutron scattering data.
arXiv:2406.05133.
External Links: LinkCited by: §I.2.
[55]A. Sapkota, L. Classen, M. B. Stone, A. T. Savici, V. O. Garlea, A. Wang, J. M. Tranquada, C. Petrovic, and I. A. Zaliznyak (2020-01)Signatures of coupling between spin waves and Dirac fermions in ${\mathrm{YbMnBi}}_{2}$.
Phys. Rev. B101 (4), pp. 041111.
External Links: Link,
DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[56]A. J. Schultz, M. R. V. Jørgensen, X. Wang, R. L. Mikkelson, D. J. Mikkelson, V. E. Lynch, P. F. Peterson, M. L. Green, and C. M. Hoffmann (2014)Integration of neutron time-of-flight single-crystal bragg peaks in reciprocal space.
J. Appl. Crystallogr.47 (3), pp. 915–921.
External Links: LinkCited by: §I.2.
[58]G. M. Sheldrick (2015)Crystal structure refinement with shelxl.
Acta Crystallogr. C71 (1), pp. 3–8.
External Links: LinkCited by: §I.2.
[59]A. K. Singh, N. Heinsdorf, A. A. Mancilla, J. Bannies, A. Maity, A. I. Kolesnikov, M. Matsuda, M. B. Stone, M. Franz, J. Gaudet, et al. (2025)Chiral spin-split magnons in the metallic altermagnet crsb.
arXiv:2511.16086.
Cited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[70]S. Toth and B. Lake (2015)Linear spin wave theory for single- incommensurate magnetic structures.
J. Phys. Condens. Matter27 (16), pp. 166002.
Cited by: §I.5.
Single crystals of -MnTe and Li-doped -MnTe were synthesized using Te self-flux method following the procedures as described in our previous work [44]. Magnetic susceptibility measurements were performed on a single-crystal Li-doped -MnTe sample with a mass of 90.4 mg, using a Quantum Design Magnetic Property Measurement System (MPMS) equipped with a superconducting quantum interference device (SQUID) magnetometer. The sample was mounted on the sample holder using GE varnish. Both field-cooled (FC) and zero-field-cooled (ZFC) measurements were conducted with the magnetic field applied along the -axis and within the -plane.
I.2 Neutron diffraction
Neutron diffraction measurements were carried out using the single crystal diffractometer TOPAZ at the Spallation Neutron Source (SNS), Oak Ridge National Laboratory (ORNL) [14]. The experiments employed a broad neutron wavelength range of 0.4–3.5 Å. Single crystals of Li-doped -MnTe ( 100 mg) and pure -MnTe ( 50 mg) were used. The samples were mounted on the ambient goniometer, and data were collected using time-of-flight (TOF) techniques, which provided wavelength-resolved Laue diffraction patterns. Diffraction datasets for both samples were obtained at 90 K and 340 K. A nitrogen cryostream was utilized for temperature control.
The data collection strategy was determined using the planning tool in NeuXtalViz [49]. A multiresolution machine learning method was employed for peak integration to obtain peak intensities in three-dimensional space [54]. Data reduction and normalization, including the Lorentz factor, neutron TOF spectrum, and detector efficiency corrections, were carried out following the procedure reported previously [56]. The reduced data were saved in SHELX HKLF2 format [58] with the neutron wavelength recorded separately for each reflection. The neutron crystal structures were solved in JANA2020 [53] and refined successfully to convergence. Crystal structures presented in the main text were visualized using the VESTA software package [47].
The momentum transfer in 3D reciprocal space is defined as where , and are Miller indices and , , with , and .
I.3 Inelastic neutron scattering
Inelastic neutron scattering (INS) measurements on undoped -MnTe and preliminary measurements on Li-doped -MnTe were performed using the fine-resolution Fermi chopper spectrometer SEQUOIA at the Spallation Neutron Source (SNS), Oak Ridge National Laboratory (ORNL) [29, 30]. For the pure -MnTe sample, a single crystal with a mass of approximately 0.46 g was mounted in the [, 0, ] scattering plane. Data were collected at K using an incident energy of meV with the high-flux chopper operating at a frequency of Hz. Measurements were performed over a total rotation range of 120 degrees about the vertical axis of the sample, with a step size of 0.5 degrees. During the experiment, the SNS accelerator operated at a beam power of 1.8 MW, and data were collected for an accumulated proton charge of approximately 0.4 Coulombs per angle step, corresponding to a counting time of roughly 5 minutes. The data obtained from SEQUOIA are presented in Figs. 2(b)–(e), 3(a)(c)(f)–(i), 4, and S2.
For the Li-doped -MnTe sample, an early-stage experiment was conducted at K using an array of co-aligned crystals with a total mass of approximately 0.4 g, also mounted in the [, 0, ] scattering plane. Several experimental configurations were employed: high-flux mode with incident energy meV at Fermi frequency Hz and meV at Hz, as well as high-resolution mode with meV at Hz.
INS measurements on Li-doped -MnTe were carried out using the TOF spectrometer MAPS at the ISIS Neutron and Muon Source, Rutherford Appleton Laboratory [23]. The sample, with a mass of approximately 2.2 g, was obtained from the same large single crystal used for the TOPAZ experiment and was mounted in the [, 0, ] scattering plane. Data were collected at K with an incident energy of meV under continuous rotation about the vertical axis of the sample. During the measurements, the accelerator operated with a synchrotron current of 152 A on target station 1. Measurements were carried out in two modes: high-resolution (HR) mode with a Fermi chopper frequency of Hz and a sample rotation range of 60 degrees, and high-flux (HF) mode with Hz and a rotation range of 120 degrees. Each mode was allocated approximately 60 hours of beamtime, evenly distributed over the full angle range. The data were reduced into .nxspe files using a step of 0.5 degrees. Data from the HF mode are presented in Figs. 2(a)(c)–(e), 3(b)(d)(f)–(i), 4, and S2, while data from the HR mode are shown in Figs. 2(c)–(e), S3, and S4.
Data reduction was performed using MANTID [3], Horace [22], and SHIVER [1].
Figure S1: Energy resolutions calculated using PyChop for the three experimental configurations used to measure the spin excitation data presented in this Letter. HF: high-flux mode; HR: high-resolution mode.
I.4 Absolute unit normalization
INS data were normalized to absolute units during the data reduction process by measuring a standard vanadium sample of known mass [72]. The vanadium measurements were performed in the same sample environment and with the same chopper and instrument settings as the sample measurements.
I.5 Linear spin wave theory (LSWT) Calculations
The spin Hamiltonian was determined from the INS spectra using linear spin wave theory (LSWT) implemented in the SpinW package [70]. For both compounds, a Heisenberg spin Hamiltonian with a single-ion anisotropy term was employed, expressed as
where indexes the Mn2+ ions, is the spin operator at site , and denotes its component. represents the exchange interaction between spins at sites and , with the summation taken over spin pairs. In this coordinate system, the -axis corresponds to the crystallographic -axis. The anisotropy term is positive for easy-plane anisotropy, as in the case of pure -MnTe, and negative for Ising-type anisotropy, as observed in Li-doped -MnTe.
To extract the exchange coupling parameters using the SpinW program,
we first constructed the lattice and magnetic structures and enabled the
relevant exchange interaction and single-ion anisotropy terms.
The lattice and magnetic structures were obtained from single-crystal neutron
diffraction results and CIF files available in the ICSD database.
The extracted spin-wave dispersion relation was then used
as input for the SpinW calculations. To determine , we performed Gaussian fits to the inelastic neutron
scattering spectra at each wave vector along high-symmetry directions.
From these fits, we extracted the peak center energy and the half-width at
half-maximum (HWHM).
For the models “Pure -MnTe center” and “Li-doped -MnTe center”
listed in Table S1, the fitted peak energies at each
along the high-symmetry directions were directly used in the SpinW
calculations. Unless otherwise specified, all spin wave calculations presented in the main text for Li-doped MnTe are based on the exchange coupling parameters obtained from “Li-doped -MnTe center” model.
For the models “Li-doped -MnTe upper” and “Li-doped
-MnTe lower,” we introduced a modification to account for
spin wave broadening. At the Brillouin zone (BZ) boundaries (i.e.,
along the M–L and L–A paths), we used and
, respectively, to represent the upper and lower
bounds of the broadened dispersion. Along all other high-symmetry directions,
we used the unmodified fitted peak energy , since the apparent broadening there is mainly
influenced by the degeneracy of optical and acoustic modes.
Table S1: Exchange coupling constants calculated using SpinW.
Sample and data
Pure -MnTe
3.8951
-0.139
0.46383
0.05519
-0.01378
0.03986
Li-doped -MnTe upper
3.6575
-0.18698
0.47811
-0.2411
N/A
Li-doped -MnTe center
3.5909
-0.22146
0.45177
-0.24515
Li-doped -MnTe lower
3.3283
-0.21053
0.47816
-0.24695
I.6 Electrical transport measurements
The longitudinal and Hall resistivities were measured using a standard five-probe configuration. Measurements were performed in a 9 T Quantum Design PPMS DynaCool system. The voltage signals were first amplified by appropriate preamplifiers (SR554) and subsequently read out using SR830 lock-in amplifiers. The sample was polished into a bar with dimensions of about . During the measurements, the magnetic field was applied along the crystallographic axis , while the electrical current flowed along the direction. All field-dependent data were symmetrized and antisymmetrized to correct for contact misalignment.
All ARPES data were taken at beamline 7.0.2 of the Advanced Light Source at Lawrence Berkeley National Lab using the deflector mode of a Scienta R4000 analyzer with an energy resolution of less than 20 meV. The samples were cleaved in-situ with a base pressure better than Torr. All measurements were performed at a temperature of 15 K using 120 eV photons with linear horizontal polarization.
II Single crystal neutron diffraction
II.1 Experiment and refinement details
Table S2: -MnTe at both temperatures. At 90 K, two magnetic structures were tested.
90 K
340 K
Space group
(No. 194)
Magnetic space group
N/A
(Å)
4.1300(5)
4.1522(5)
(Å)
6.6530(11)
6.7162(11)
(Å3)
98.28(2)
100.28(2)
Index range
-8<<9
-8<<9
-14<<14
-9<<8
-8<<9
-14<<14
Reflections collected
((obs)/(all))
2137/2154
1460/1468
indices
( >3)
=3.57
=10.89
=3.47
= 9.97
=3.13
=7.69
R indices (all data)
=3.63
=10.92
= 3.52
=10.01
=3.16
=7.70
Goodness-of-fit
2.84
2.16
1.85
Moment size ()
4.35(3)
4.30(3)
N/A
For both -MnTe and Li-doped
-MnTe, we find no evidence of any other impurity phases, clearly different from powder samples of Li-doped -MnTe [73].
For -MnTe, we also evaluated the possibility of intersite mixing, specifically the occupation of Te sites by Mn atoms, using the neutron diffraction dataset collected at 340 K. However, this model did not yield an improved refinement compared to the current model, which assumes that defects are restricted to Te sites and that Mn atoms exclusively occupy their designated lattice positions.
To determine the magnetic structure, two models were tested using the dataset obtained at 90 K. The first model adopts the magnetic space group , where the magnetic moments align along the crystallographic axis. The second model corresponds to the magnetic space group , in which the moments are oriented along the reciprocal axis. Although the model yielded a slightly better refinement result, the model is more consistent with our previous neutron diffraction results under strain-induced detwinning of magnetic domains [44].
Table S3: Li-doped -MnTe at 340 K. Models placing Li at Wyckoff site , site , site , or uncorrelated interstitial sites were tested.
Li site
4
2
2
Uncorrelated
Space group
(No. 194)
4.1532(3)
6.7190(8)
100.369(16)
Index range
-99
-78
-1314
Reflections collected
((obs)/(all))
1980/2002
indices
[…]
= 3.52
= 11.74
=3.52
=11.76
=3.51
= 11.76
=3.52
=11.76
indices (all data)
= 3.60
= 11.76
=3.60
=11.78
= 3.59
=11.78
=3.60
=11.78
Goodness-of-fit
2.99
2.99
2.99
3.00
We used the 340 K diffraction dataset to investigate the possible crystallographic location of Li atoms. As a first scenario, we considered the possibility that Li atoms occupy uncorrelated interstitial sites. In this case, the Li atoms would not contribute coherent scattering, and therefore the Bragg peak intensities would remain unaffected. Accordingly, Li atoms were not included in the refinement model for this scenario.
We then tested alternative structural models incorporating Li atoms at
different crystallographic sites. Stable refinements were obtained when
Li was placed at the , , and positions (the latter coinciding
with the Te site). However, when Li atoms were assigned to the site,
the refined concentration was significantly lower than the nominal doping
level of 5%, as shown in Table S10. A similar refinement using the 90 K dataset yielded negative Li occupancies
at the site when the atomic displacement parameters (ADPs) were not
constrained, which is physically implausible. Placing Li at the site
resulted in a refined concentration of 1.7%, still much smaller than
the nominal doping level. Moreover, refinement of the 340 K dataset in this
model produced unphysical negative anisotropic ADPs. Therefore, we conclude that neither the nor the site provides a
physically reasonable model to account for the observed neutron diffraction
intensities.
In contrast, assigning Li atoms to the 2 site resulted in stable refinements at both temperatures, with a consistent refined concentration of approximately 3%, suggesting that the 2 site is a more plausible position for Li incorporation. Nevertheless, we note that the inclusion of Li atoms at either site did not significantly improve the overall refinement quality (Table S3). Therefore, the possibility that Li atoms reside at random interstitial positions cannot be ruled out.
Table S4: Li-doped -MnTe at 90 K.
Magnetic space group
Space group
(No. 194)
4.130(2)
6.666(2)
98.48(7)
Index range
-9<<9
-7<<8
-13<<14
Reflections collected
((obs)/(all))
3215/3261
indices
( >3)
=4.65
=16.04
=4.71
= 16.14
indices (all data)
=4.71
=16.06
= 4.76
=16.16
Goodness-of-fit
4.27
4.30
Moment size ()
4.29(3)
4.30(10)
In accordance with previous powder neutron diffraction studies [48], we applied the orthorhombic magnetic space group to refine the magnetic structure of Li-doped -MnTe. This model also yielded a vanishingly small in-plane magnetic moment at the given doping level. In addition, we tested a hexagonal magnetic space group with higher symmetry, in which the magnetic structure is strictly antiferromagnetic and the magnetic moments are constrained to lie along the crystallographic axis. The refinement results for both models are summarized in Table S4.
II.2 Fractional atomic coordinates and displacement parameter (Å2) determined from neutron diffraction
Table S5: MnTe, 340 K
Atom
Label
Occ.
Site
Sym.
Mn
Mn1
0
0
0
1.000
0.017
2
Te
Te1
1/3
2/3
1/4
0.967(2)
0.010
2
Table S6: MnTe, 90 K
Atom
Label
Occ.
Site
Sym.
Mn
Mn1
0
0
0
1.000
0.004
2
Te
Te1
1/3
2/3
1/4
0.957(2)
0.002
2
Table S7: Li-doped -MnTe, without Li atoms
Atom
Label
Occ.
Site
Sym.
Mn
Mn1
0
0
0
1.000
0.017
2
Te
Te1
1/3
2/3
1/4
0.961(3)
0.010
2
Table S8: Li-doped -MnTe, 340 K with the assumption at Li atoms sit at the Te site
Atom
Label
Occ.
Site
Sym.
Mn
Mn1
0
0
0
1.000
0.017
2
Te
Te1
1/3
2/3
1/4
0.970(1)
0.010
2
Li
Li1
1/3
2/3
1/4
0.030(1)
0.010
2
Table S9: Li-doped -MnTe, 340 K with the assumption at Li atoms sit at the 2 site
Atom
Label
Occ.
Site
Sym.
Mn
Mn1
0
0
0
1.000
0.018
2
Te
Te1
1/3
2/3
1/4
0.962(4)
0.010
2
Li
Li1
1/3
2/3
3/4
0.007(16)
0.041
2
Table S10: Li-doped -MnTe, 340 K with the assumption at Li atoms sit at the 4 site
Atom
Label
Occ.
Site
Sym.
Mn
Mn1
0
0
0
1.000
0.018
2
Te
Te1
1/3
2/3
1/4
0.959(3)
0.010
2
Li
Li1
1/3
2/3
0.898(14)
0.009(6)
0.02
4
Table S11: Li-doped -MnTe, 90 K with the assumption at Li atoms sit at the Te site
Atom
Label
Occ.
Site
Sym.
Mn
Mn1
0
0
0
1.000
0.004
2
Te
Te1
1/3
2/3
1/4
0.969(2)
0.002
2
Li
Li1
1/3
2/3
1/4
0.031(2)
0.002
2
III INS data for Li-doped -MnTe taken using the HR mode
Figure S2: Spin waves in Li-doped -MnTe along high-symmetry directions in high-resolution mode.
Figure S3: (a) INS spectra of Li-doped -MnTe along the and (b) directions in high-resolution mode.
IV Energy-dependence of magnon lifetime in Li-doped -MnTe
Figure S4: (a) Energy-dependence of the full width at half-maximum (FWHM) for Li-doped -MnTe in high-flux (HF) mode (filled symbols) and high-resolution (HR) mode (open symbols). Dashed line serves as a guide to the eye marking the onset of increased linewidth above meV. (b) Differences in FWHM (FWHM) between Li-doped -MnTe and -MnTe along various high-symmetry directions in the nodal planes of the BZ. The top panel utilizes HF data for both Li-doped -MnTe and -MnTe, where the resolution is comparable between Li-doped -MnTe and -MnTe measurements. The bottom panel uses HR data for Li-doped -MnTe and HF data for -MnTe, meaning the resolutions are different for each.
FWHM values were subtracted at identical for both samples, and the corresponding energy of the Li-doped -MnTe dataset () is used as the -axis value for each point.
V Additional transport data
Figure S5: Longitudinal and transverse resistivity and of pure and Li-doped -MnTe.
VI Local moment sum rule
The total moment sum rule requires , where is the static ordered moment contribution (as obtained from the refined magnetic structure using neutron diffraction) and
is the dynamic contribution where is the -averaged local dynamic susceptibility, , and is the Boltzmann constant [16].
is directly related to the measured quantity from INS experiments, and can be integrated over a full BZ to yield the energy dependence of the local dynamic susceptibility . The -averaged local dynamic susceptibility is given by
where the subscript BZ denotes -integration over a full Brillouin zone. The relationship between and the neutron scattering intensity measured during the experiment is given by
where is assumed to be in absolute units of b sr-1 f.u.-1 meV-1, 1 b = 10-24 cm2 is the unit “barn” for the neutron scattering cross section, is the magnetic form factor, and is the Debye-Waller factor [72].
-integration of the total spectral weight requires enough coverage to capture a full Brillouin zone, or, at least, enough coverage to reconstruct the full Brillouin zone from a smaller, symmetric portion of it. Because the accessible range of momentum transfer at a given energy transfer is restricted by the kinematic condition, coverage gradually shifts to higher at higher energy transfers. As a result, the coverage at smaller [e.g., ] is better at low energies, and the coverage at larger [e.g., ] is better at high energies. In order to capture the total spectral weight over the full energy range of the spin waves, was used for lower energy transfers, and was used for higher energy transfers, with the specific range informed by the coverage of the respective instrument for MAPS and SEQUOIA datasets.
VII Momentum- and energy-integration ranges for the data
In Figures 2, S2, and S3, the momentum integration directions and the plotting direction (if applicable) are mutually perpendicular. The integration ranges were chosen to ensure that the integrated volume in reciprocal space (in units of Å-1) is consistent across different directions throughout each figure. Integration ranges for other figures are provided in Table S14. For datasets where altermagnetic splitting is expected, such small integration ranges may be necessary to resolve the splitting clearly.
Table S12: Momentum-integration ranges for Fig. 2 (a)-(b) and S2
Path
Momentum 1 (r.l.u)
Momentum 2 (r.l.u)
-M
[, , 0], 0.0578
0, 0, ], 0.093
M-L
[, 0, 0], 0.05
[0, 0, ], 0.093
L-A
[, , 0], 0.0578
[0, 0, ], 0.093
-A
[, 0, 0], 0.05
[0, 0, ], 0.093
-H
[, , 0], 0.05
[, , ], 0.012
-K-M
[, , 0], 0.05
[0, 0, ], 0.093
Table S13: Momentum-integration ranges for Fig. 2(c)-(e)
Path
Momentum 1 (r.l.u)
Momentum 2 (r.l.u)
Momentum 3 (r.l.u)
K-
[, , 0], 0.04
[0, , 0], 0.03
[0, 0, ], 0.06
-M
[, 0, 0], 0.03
[, , 0], 0.03
[0, 0, ], 0.06
M-L
[0, 0, ], 0.06
[, , 0], 0.03
[0, , 0], 0.03
L-A
[, 0, 0], 0.03
[, , 0], 0.03
[0, 0, ], 0.06
-A
[0, 0, ], 0.06
[, , 0], 0.03
[0, , 0], 0.03
-H
[, , 0], 0.04
[0, , 0], 0.03
[, , ], 0.009
Table S14: Momentum- and energy-integration ranges for other figures